OpenStudy (anonymous):

give me the ordered pairs please :) 1. y = –x + 2 y = 3x – 1 2. –4x + 3y = –12 –2x + 3y = –18

OpenStudy (anonymous):

-10x+10y=20 -15x+5y=-5